OpenStudy (anonymous):

if a stone is thrown into the sea, the distance s(t) it travels after t seconds is s(t) = aln( cosh bt ), where a and b are positive constants. find the velocity v and evaluate v(30)..?.. *how do i differentiate s(t) = aln( cosh bt )..?

OpenStudy (anonymous):

Use the chain rule.

OpenStudy (anonymous):

\[ u(t) = \cosh(bt)\\ s(t) = a\ln(u(t)) \]

OpenStudy (anonymous):

\[ du = \sinh(bt)dt\\ ds = \frac{a\;du}{u} \]

OpenStudy (anonymous):

Now, we can substitute for \(du\) and \(u\): \[ ds = \frac{ab\sinh(bt)\;dt}{\cosh(bt)} = ab\tanh(bt)dt \]And so \[ s'(t) = \frac{ds}{dt} = ab\tanh(bt) \]

OpenStudy (anonymous):

Should have said \[ du = b\sinh(bt)dt \]up above. It was a typo.

OpenStudy (anonymous):

then how do i evaluate v(30)..?.. did i simply substitute it..?..

OpenStudy (anonymous):

Well, \(v(t) = s'(t)\). So \(v(30) = s'(30)\).

OpenStudy (anonymous):

so v(30) = abtanh(b(30))..? am i right..?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

is it possible to evaluate it further.,. i mean making v(30) = a number like 34, 50 or any ather number..?

OpenStudy (anonymous):

No matter what you do, the answer would likely still be a function of \(a\) and \(b\) since they are variables which have not been specified.

OpenStudy (anonymous):

I think it is simplified enough as is. I'm not sure what hyperbolic identities would be used to simplify further.

OpenStudy (anonymous):

ok thanks.... by the way how do i send you a medal..?

OpenStudy (anonymous):

"Best Answer" I think.

OpenStudy (anonymous):

Click on "Best Response"

OpenStudy (anonymous):

Blue button in top right corner.

OpenStudy (anonymous):

^

OpenStudy (anonymous):

do you receive a medal from me... i just click the *best response* button..?

OpenStudy (anonymous):

yeah